Question NUMBER 8 only please Verify the initial case. State the induction hypothesis. Perform the induction. See Example 5.2.1. 6. (6 pts) Prove by mathem
Answer #1

Basis step for n= os for n=1, that s bo= qo-4 = 0 bi=91-41=5. divides both bo & 6. So, it is clear Inductive step: Assue that